OpenStudy (aravindg):

Doubts:

OpenStudy (aravindg):

4 1 ohm resistances of tolerance 5% each are connected in series. Find the tolerance of equivalent resistance?

OpenStudy (aravindg):

@Vincent-Lyon.Fr

OpenStudy (atlas):

for resitances in series: Tolerance = \[\frac{ \sum_{}^{}RT }{ \sum_{}^{}R }\]

OpenStudy (vincent-lyon.fr):

Sorry, I'm not an expert in this field, but I think \(R_{eq}=R_1+R_2+R_3\) \(\dfrac{\Delta R_{eq}}{R_{eq}}=\dfrac{\Delta (R_1+R_2+R_3)}{R_1+R_2+R_3}=\dfrac{3\Delta R_1}{3R_1}=\dfrac{\Delta R_1}{R_1}=5\)%

OpenStudy (kropot72):

An interesting approach to this problem can be found through a statistical analysis which assumes a normal distribution of the values (although in practice the distribution might not be normal). Such an analysis can be found here: http://paulorenato.com/joomla/index.php?option=com_content&view=article&id=109&Itemid=4 Given n resistors of mean value \[E(R _{i})=\mu\] and tolerance \[T _{i}\]

OpenStudy (kropot72):

the tolerance of the resulting resistor Tt is \[\frac{T _{i}}{\sqrt{n}}\]

OpenStudy (kropot72):

So taking the values in the question we get the initial tolerance of 5% for each resistor reduced to \[\frac{5}{\sqrt{4}}=2.5\ percent\] for the resulting total resistance.

OpenStudy (atlas):

So at the worst (with the most poorly designed resistors within the toloerance limits) the tolerance will be 5% though NORMALIZED tolerance is 2.5%

OpenStudy (aravindg):

The options given were 5%,10%,15%,20%. I guess I will go with Vince's method.

OpenStudy (kropot72):

If the values followed a continuous uniform (rectangular) distribution the tolerance of the equivalent resistance would indeed be 5%. However such a distribution would clearly show that the manufacturer selected the resistances to meet the specified tolerance rather than manufactured them to meet a specific mean value by means of a well controlled process.

OpenStudy (aravindg):

Thanks @kropot72 :)

OpenStudy (kropot72):

You're most welcome :)